OpenStudy (anonymous):

A population of salamanders doubles every month. There are 15 salamanders in the backyard lake. If nothing cuts their numbers, how many will there be in twelve months? Let months be t, and “the population of salamanders at time t” be S(t).

jigglypuff314 (jigglypuff314):

the equation would be S(t)=2t+15 you would plug 12 in so 2(12)+15= (answer)

OpenStudy (anonymous):

I'm not so sure about that one. If i do that then i get 39 but it clearly tells me that they double every month. I know what the total should be just by adding it up manually, just trying to figure out the equation for it.

OpenStudy (ranga):

This is a geometric series: 15, 30, 60, 120, 240, .... The nth term of a geometric series: a, ar, ar^2, ar^3 is: nth term = ar^(n - 1) where a = 15, r = 2 nth term = (15)2^(n - 1) They want the formula using t for n: s(t) = (15)2^(t - 1) Put t = 12 in the ab0ove equation and that should be your answer.

OpenStudy (ranga):

30,720?

OpenStudy (anonymous):

Yes that's the total i got by simply doubling the qty every month. Thanks for explaining the formula for this problem.

OpenStudy (ranga):

Glad to help.